# Enhancing Training Efficiency Using Packing with Flash Attention

### v3

@@ Abstract
Padding is often used in tuning LLM models by adding special tokens to shorter training examples to match the length of the longest sequence in each batch. While this ensures uniformity for batch processing, it introduces inefficiencies by including irrelevant padding tokens in the computation and wastes GPU resources. On the other hand, the Hugging Face SFT trainer offers the option to use packing to combine multiple training examples up to the maximum sequence length. This allows for maximal utilization of GPU resources. However, without proper masking of each packed training example, attention will not be computed correctly when using SFT trainer. We enable and then analyse packing and Flash Attention with proper attention masking of each example and show the benefits of this training paradigm.
@@ 2 Related Work
There are a few ways that sample packing can be enabled with Flash Attention 2 [9] and proper attention masking of each example. The Flash Attention repository itself offers a way to pack while enabling proper masking of examples with Flash Attention. See Flash Attention Closed Issue 654 [1]. Others have proposed padding-free transformers, such as [6, 7]. The padding-free transformer methods require substantial and intrusive changes however to Hugging Face transformers library. As such, these methods have seen less uptake by the community and are not currently available in the Hugging Face library. For Hugging Face SFT Trainer users, it is desirable to have a readily available solution without requiring going outside the library.
@@ 5 Conclusions
We provide a lightweight solution for packing with PositionIDs that can be readily integrated into Hugging Face transformers library and demonstrate its benefits.

### v4

@@ Abstract
Padding is often used in tuning LLM models by adding special tokens to shorter training examples to match the length of the longest sequence in each batch. While this ensures uniformity for batch processing, it introduces inefficiencies by including irrelevant padding tokens in the computation and wastes GPU resources. Hugging Face SFT trainer has always offered the option to use packing to combine multiple training examples, allowing for maximal utilization of GPU resources. However, up till now, it did not offer proper masking of each packed training example. This capability has now been added to Hugging Face Transformers 4.43. We analyse this new feature and show the benefits across different variations of packing.
@@ 2 Related Work
There are a few ways that sample packing can be enabled with Flash Attention 2 [9] and proper attention masking of each example. The Flash Attention repository itself offers a way to pack while enabling proper masking of examples with Flash Attention. See Flash Attention Closed Issue 654 [1]. Others have proposed padding-free transformers, such as [6, 7]. The padding-free transformer methods require substantial and intrusive changes however to Hugging Face transformers library. As such, these methods have seen less uptake by the community and are not currently available in the Hugging Face library. For Hugging Face SFT Trainer users, it is desirable to have a readily available solution without requiring going outside the library. This feature is now available in Transformers version 4.43.
@@ 3.1 Online Mini-batch Collating
This capability has been added to Hugging Face trainer from Transformers 4.43, leveraging the new DataCollatorWithFlattening.
@@ 5 Conclusions
We provide a lightweight solution for packing with PositionIDs that has been integrated into Hugging Face Transformers library version 4.43 and we demonstrate its benefits.